# Contribution of remote bands to orbital magnetization in twisted bilayer graphene

### 04 Mar 2026 v1

@@ Abstract
Motivated by recent theoretical and experimental works on orbital magnetization $M_{\mathrm{orb}}$ for the interacting system, we develop a gauge-invariant framework to compute $M_{\mathrm{orb}}$ for correlated phases of magic-angle twisted bilayer graphene within self-consistent Hartree-Fock approximation. Based on the projector formulation of the theory of orbital magnetization, we evaluate both $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ directly from the Hartree–Fock Hamiltonian. We demonstrate that, in contrast to topological invariants such as the Chern number, both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ obtain substantial contributions from remote bands and thus require careful convergence with respect to the number of included remote bands. Applying this approach to correlated phases at integer fillings, we obtain converged $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ for time reversal symmetry broken Chern insulating states at $\nu=\pm 3$ and for competing correlated phases at other integer fillings. Our results establish a systematic and controlled approach for evaluating orbital magnetization in correlated moiré systems and clarify the crucial role of remote bands in determining their magnetic response.
@@ I Introduction
The recent discovery of the correlated insulating states Cao et al. (2018a); Sharpe et al. (2019); Lu et al. (2019); Jiang et al. (2019); Kerelsky et al. (2019); Xie et al. (2019); Choi et al. (2019); Serlin et al. (2020); Yankowitz et al. (2019) and unconventional superconductivity Cao et al. (2018b); Lu et al. (2019); Yankowitz et al. (2019); Codecido et al. (2019) in twisted bilayer graphene (TBG) has stimulated tremendous research efforts in moiré systems. Due to the extreme flatness of the lowest energy bands in TBG near the magic angle Bistritzer and MacDonald (2011) and their nontrivial topology Song et al. (2019, 2021); Ahn et al. (2019); Tarnopolsky et al. (2019); Liu et al. (2019a); Po et al. (2019), electronic Coulomb interactions are believed to drive a rich and robust set of strongly correlated and topological phases. Theoretical studies have predicted several strongly correlated and topologically distinct states at integer fillings  Kang and Vafek (2019, 2020); Zhang et al. (2019, 2020); Xu et al. (2018); Xie and MacDonald (2020); Liu et al. (2018); Huang et al. (2019); Liu and Dai (2021a); Liu et al. (2021); Bultinck et al. (2020a, b); Wu and Das Sarma (2020); Chen et al. (2021); Lu et al. (2022); Da Liao et al. (2021); Bernevig et al. (2021); Lian et al. (2021); Xie et al. (2021); Hejazi et al. (2021); Sboychakov et al. (2019); Rademaker et al. (2019); Guinea and Walet (2018); Seo et al. (2019). Among these states, the time-reversal symmetry breaking (TRSB) phases can become Chern insulators, exhibiting the quantum anomalous Hall (QAH) effect, which have been observed experimentally Sharpe et al. (2019); Serlin et al. (2020); Wu et al. (2021); Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022). The ferromagnetic QAH states at $\nu=3$ filling was discovered to show strong magnetic response to external DC currents Sharpe et al. (2019); Serlin et al. (2020), which can be understood from the large orbital magnetization of the TBG system as a result of TRSB He et al. (2020); Su and Lin (2020); Huang et al. (2021). Specifically, this strong orbital effect has been studied in moiré systems where the time-reversal symmetry is explicitly broken Lee et al. (2019); Zhang et al. (2022), and in the $C_{2}\mathcal{T}$ symmetric moiré system such as TBG in the presence of additional $C_{2}\mathcal{T}$ symmetry breaking potential He et al. (2020); Su and Lin (2020).
@@
Moreover, the transport and scanning tunneling microscopy (STM) experiments revealed that weak to moderate out-of-plane magnetic fields can turn the ground states at $\nu=\pm 1,\pm 2,\pm 3$ into strong-correlation induced Chern insulators in magic-angle TBG Das et al. (2021); Nuckolls et al. (2020). The associated orbital magnetization plays a central role in these phenomena, which determines the coupling to external magnetic fields. Understanding orbital magnetization in correlated moiré systems is therefore essential for elucidating the microscopic origin of Chern insulators and field-induced quantum phase transitions in TBG Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022).
@@
Theoretically, the Berry curvature and orbital magnetic moments of moiré bands have been extensively analyzed within the non-interacting Bistritzer–MacDonald (BM) continuum model Liu et al. (2019b); He et al. (2020); Liu and Dai (2021b); Zhu et al. (2020). These studies established that the flat bands of TBG can host nontrivial topology and substantial orbital magnetization per unit cell largely arising from sizable Berry curvature. However, the flat-band nature of magic-angle TBG implies that electron–electron interactions are comparable to or even larger than the bandwidth, and self-consistent Hartree–Fock (HF) calculations have demonstrated that interactions tremendously reconstruct the band structure and stabilize correlated Chern insulating states at integer fillings Kang and Vafek (2019, 2020); Zhang et al. (2019, 2020); Xu et al. (2018); Xie and MacDonald (2020); Liu et al. (2018); Huang et al. (2019); Liu and Dai (2021a); Liu et al. (2021); Bultinck et al. (2020a, b); Wu and Das Sarma (2020); Chen et al. (2021); Lu et al. (2022); Da Liao et al. (2021); Bernevig et al. (2021); Lian et al. (2021); Xie et al. (2021); Hejazi et al. (2021); Sboychakov et al. (2019); Rademaker et al. (2019); Guinea and Walet (2018); Seo et al. (2019). Despite this progress, a systematic calculation of the bulk orbital magnetization in such interaction-driven states remains lacking. In particular, it is not clear how interaction-induced band reconstruction, valley polarization, and remote-band contributions quantitatively modify the orbital magnetization from the non-interacting predictions, nor how these effects relate to experimentally observed field-induced switching between competing correlated phases.
@@
In this work, we adopt the recently-developed method Kang et al. (2025); Zhu and Huang (2025); Liu et al. (2025); Ye (2026); Chen and Song (2026) and develop a gauge-invariant framework Ceresoli et al. (2006) for computing orbital magnetization within self-consistent Hartree–Fock theory and apply it to correlated insulating states of magic-angle TBG. Starting from the BM continuum model with Coulomb interactions projected onto the relevant bands, we determine symmetry-broken ground states at integer fillings and evaluate the orbital magnetization using the modern theory expressed in terms of projection operators. This formulation avoids gauge ambiguities and enables a controlled treatment of the large number of remote bands inherent to the continuum description. We demonstrate that, unlike topological invariants such as the Chern number, the orbital magnetization exhibits a pronounced sensitivity to remote-band contributions and requires careful convergence. Applying this framework, we compute both the total orbital magnetization $M_{orb}$ and the self-rotation (local circulation) component $m_{SR}$ relevant for magneto-optical measurements such as magnetic circular dichroism (MCD). The calculated orbital magnetization respects the particle-hole symmetry that relates the fillings of $\nu$ and $\nu$ Bernevig et al. (2021) and reproduces the characteristic linear dependence of magnetization on chemical potential within the insulating gap, with slope proportional to the Chern number. We further analyze the sign structure of the magnetization and its relation to changes in Hall conductivity. At $\nu=\pm 1$, we compare competing Chern and intervalley-coherent states and show that Chern insulating phases possess substantially larger orbital magnetization, providing a microscopic explanation for their stabilization under external magnetic fields. Our results establish a systematic approach for evaluating orbital magnetization in correlated moiré systems and elucidate the important role of interactions as well as the critical contribution of remote bands in shaping orbital ferromagnetism in twisted bilayer graphene.
@@ II Non-interacting model
where $\hat{{\mathbf{k}}}=-i\partial_{{\mathbf{r}}}$, $\tau=\pm$ is the valley index, and $\boldsymbol{\sigma}^{\tau}=(\tau\sigma_{x},\sigma_{y})$ the Pauli matrices acting in the A,B sublattice space. In Eq. (1), ${\mathbf{K}}^{\tau}_{1}$ and ${\mathbf{K}}^{\tau}_{2}$ are the Dirac points of valley $\tau$ in the bottom and top layers that are twisted by angles $\mp\frac{\theta}{2}$, which are equivalent to $\tau{\mathbf{K}}_{1}$ and $\tau{\mathbf{K}}_{2}$ in the mBZ as shown in Fig. 1(c). The interlayer tunneling between the Dirac states in the two layers is described by the matrix
@@
where $u_{0}$ and $u_{1}$ are the intra- and inter-sublattice interlayer tunneling amplitudes. We adopt realistic parameters for the continuum model: $\hbar v_{F}/a_{0}=2.365$ eV, inter-sublattice interlayer coupling $u_{1}=0.11$ eV, and intra-sublattice interlayer coupling $u_{0}=0.06$ eV. The twist angle is fixed at the magic value $\theta$=1.086, at which the low-energy moiré bands become perfectly flat in the limit $u_{0}\to 0$ Tarnopolsky et al. (2019). The low-energy flat bands exhibit a bandwidth of approximately 2 meV and are well separated from the higher-energy remote bands by an energy gap of about 50 meV as shown in Fig. 1. The eigenstates of the continuum Hamiltonian $H_{BM}^{\tau}$ have the Bloch wavefucntion form
@@
where $X=\{A_{1},B_{1},A_{2},B_{2}\}$ is the layer and sublattice index with the eigen-band energy $\epsilon_{m{\mathbf{k}}\tau}$ and $\mathbf{G}$} is the general reciprocal lattice vector in the mBZ. Here, $m$ and $\tau$ are the band and valley indices, and the spin index $s$ is omitted since the Hamiltonian is spin independent.
Figure 1: (a) Low-energy band structure of non-interacting TBG calculated using the BM model. The black solid and red dashed curves correspond to the $K$’ and $K$ valleys, respectively. (b) Zoomed-in view of the isolated flat bands near charge neutrality, highlighting their separation from the higher-energy remote bands. (c) Schematic illustration of Brillouin zone folding in TBG. The small hexagon denotes the moiré Brillouin zone (mBZ), reciprocal to the moiré superlattice. The cyan arrows indicate the momentum path in the enlarged moiré Brillouin zone along which the band dispersions in panels (a) and (b) are plotted.
@@ III Interacting case
The full Hamiltonian $H=H_{0}+H_{int}$ contains the non-interacting part
$H_{0}=\sum_{\tau}H_{BM}^{\tau}=\sum_{m{\mathbf{k}}s\tau}\epsilon_{m{\mathbf{k}}\tau}d_{m{\mathbf{k}}s\tau}^{\dagger}d_{m{\mathbf{k}}s\tau},$ (4)
where $d^{\dagger}_{m{\mathbf{k}}s\tau}$ and $d_{m{\mathbf{k}}s\tau}$ are the creation and annihilation operators in the eigen-band basis. To incorporate interactions, we project the Coulomb interaction onto the two lowest-energy flat bands per spin and per valley, and perform self-consistent Hartree–Fock calculations directly in momentum space. This approach avoids the real-space Wannier obstruction associated with topological flat bands Po et al. (2019, 2018a, 2018b); Ahn et al. (2019); Song et al. (2021). Further technical details of the Hartree–Fock procedure can be found in the Appendix of Ref.Zhang et al., 2020. The interacting part of the Hamiltonian is written asBernevig et al. (2021); Kang et al. (2021)
$H_{int}=\frac{1}{2S}\sum_{{\mathbf{q}}\neq 0}V({\mathbf{q}})\delta\rho_{{\mathbf{q}}}\delta\rho_{-{\mathbf{q}}}\ ,$ (5)
where S is the total area and $\delta\rho_{{\mathbf{q}}}=\rho_{{\mathbf{q}}}-\bar{\rho}_{{\mathbf{q}}}$ is the density operator relative to charge neutrality of the decoupled bilayer graphene in momentum space. We note that this interaction form is somewhat different from the normal ordered interactions supplemented by double-counting subtraction schemes used in previous Hartree-Fock (HF) studies Bultinck et al. (2020b); Liu et al. (2021); Zhang et al. (2020). It is believed that the Coulomb interaction expressed in Eq. 5 has the advantage of partially accounting for the renormalization and screening effects of the remote bands on the lowest flat bands Kang et al. (2021); Bernevig et al. (2021). The density operator can be expanded in the eigen-basis of $H_{0}$,
$\rho_{{\mathbf{q}}}=\sum_{s,\tau}\sum_{{\mathbf{k}}\in\text{mBZ},m_{1},m_{2}}\lambda_{m_{1},m_{2};\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{q}})d^{\dagger}_{{\mathbf{k}},m_{1},s,\tau}d_{{\mathbf{k}}+{\mathbf{q}},m_{2},s,\tau}\,$ (6)
with the quantum averaged background density
$\bar{\rho}_{{\mathbf{q}}}=\sum_{{\mathbf{G}},\tau,m}\delta_{{\mathbf{q}},{\mathbf{G}}}\sum_{{\mathbf{k}}\in\text{mBZ}}\lambda_{m,m;\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{G}})\ ,$ (7)
where $\lambda$ is the form factor defined in terms of the Bloch wavefunction as
$\begin{split}\lambda_{m_{1}m_{2};\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{q}})&=\left<u_{m_{1},\tau}({\mathbf{k}})|u_{m_{2},\tau}({\mathbf{k}}+{\mathbf{q}})\right>\\ &=\sum_{{\mathbf{G}}^{{}^{\prime}},X}u^{*}_{m_{1},\tau;{\mathbf{G}}^{{}^{\prime}},X}({\mathbf{k}})u_{m_{2},\tau;{\mathbf{G}}^{{}^{\prime}},X}({\mathbf{k}}+{\mathbf{q}}).\end{split}$ (8)
Here, $\mathbf{k}$} is restricted in the mBZ while $\mathbf{q}$} can extend to the original Brillouin zone of the monolayer graphene. The operator $d^{\dagger}$ obeys the periodic condition $d^{\dagger}_{m{\mathbf{k}}s\tau}=d^{\dagger}_{m{\mathbf{k}}+{\mathbf{G}}s\tau}$, which is equivalent to the Bloch wave function obeying the condition $u_{m,\tau;{\mathbf{G}},X}({\mathbf{k}}+{\mathbf{G}}_{0})=u_{m,\tau;{\mathbf{G}}+{\mathbf{G}}_{0},X}({\mathbf{k}})$. We take $V({\mathbf{q}})$ as the single-gate-screened Coulomb potential Bultinck et al. (2020b)
$V({\mathbf{q}})=\frac{e^{2}}{2\varepsilon\varepsilon_{0}q}(1-e^{-2qd_{s}})$ (9)
In this paper, we take the dielectric constant $\varepsilon=7$ and the gate distance $d_{s}=40$ nm.
@@
It is well established that electron–electron Coulomb interactions in twisted bilayer graphene can drive spontaneous symmetry breaking, giving rise to a series of correlated insulating states at commensurate integer fillings. Among these, states that break the combined $C_{2}\mathcal{T}$ symmetry are energetically competitive. In particular, at fillings $\nu=\pm 3$, the $C_{2}\mathcal{T}$-broken Chern insulating states can emerge as ground states even in the absence of an explicit staggered sublattice potential. In this work, we systematically identify the correlated insulating states at all integer fillings within the Hartree–Fock approximation and compute the orbital magnetization of these interaction-driven phases.

### 16 Apr 2026 v2

@@ Abstract
Motivated by recent theoretical and experimental studies of orbital magnetization $M_{\mathrm{orb}}$ in interacting systems, we develop a gauge-invariant framework to compute $M_{\mathrm{orb}}$ for correlated phases of magic-angle twisted bilayer graphene within a self-consistent Hartree–Fock approximation. Based on the projector formulation of the theory of orbital magnetization, we evaluate both $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ directly from the Hartree–Fock Hamiltonian. We demonstrate that, in contrast to topological invariants such as the Chern number, both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ receive substantial contributions from remote bands and thus require careful convergence with respect to the number of included remote bands. Applying this approach to correlated phases at integer fillings, we obtain converged $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ for time-reversal-symmetry-broken Chern insulating states at $\nu=\pm 3$ and for competing correlated phases at other integer fillings. Our results establish a systematic and controlled approach for evaluating orbital magnetization in correlated moiré systems and clarify the crucial role of remote bands in determining their magnetic response.
@@ I Introduction
The discovery of correlated insulating states Cao et al. (2018a); Jiang et al. (2019); Kerelsky et al. (2019); Xie et al. (2019); Choi et al. (2019); Xie et al. (2021a); Sharpe et al. (2019); Serlin et al. (2020); Wu et al. (2021); Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022); Lu et al. (2019); Yankowitz et al. (2019) and unconventional superconductivity Cao et al. (2018b); Lu et al. (2019); Yankowitz et al. (2019); Codecido et al. (2019) in twisted bilayer graphene (TBG) has stimulated intense interest in correlated moiré systems. Near the magic angle, the isolated low-energy bands are extremely flat Bistritzer and MacDonald (2011) and possess nontrivial topology Tarnopolsky et al. (2019); Liu et al. (2019a); Song et al. (2019, 2021); Ahn et al. (2019); Po et al. (2019, 2018a, 2018b), so electron-electron interactions are expected to drive a rich variety of symmetry-broken and topological phases. Indeed, theoretical studies have predicted several strongly correlated states at integer fillings Kang and Vafek (2019, 2020); Zhang et al. (2019); Xu et al. (2018); Xie and MacDonald (2020); Liu et al. (2018); Huang et al. (2019); Liu and Dai (2021a); Liu et al. (2021); Bultinck et al. (2020a, b); Wu and Das Sarma (2020); Chen et al. (2021); Lu et al. (2022); Da Liao et al. (2021); Lian et al. (2021); Xie et al. (2021b); Hejazi et al. (2021); Sboychakov et al. (2019); Rademaker et al. (2019); Guinea and Walet (2018); Seo et al. (2019); Bernevig et al. (2021); Kang et al. (2021); Zhang et al. (2020); Wang and Vafek (2024, 2022). Among these phases, time-reversal symmetry breaking (TRSB) states are particularly interesting because they can become Chern insulators and exhibit integer or fractional quantum anomalous Hall (QAH) effects, as observed experimentally in TBG Sharpe et al. (2019); Serlin et al. (2020); Wu et al. (2021); Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022); Xie et al. (2021a) and other moiré materials Chen et al. (2020); Li et al. (2021); Xu et al. (2023); Park et al. (2023); Lu et al. (2024). In particular, the ferromagnetic QAH state at $\nu=3$ was found to exhibit a strong magnetic response to external dc currents Sharpe et al. (2019); Serlin et al. (2020), pointing to a large orbital magnetization associated with TRSB He et al. (2020); Su and Lin (2020); Huang et al. (2021); Lee et al. (2019); Zhang et al. (2022). Orbital magnetization in TBG has also been probed directly by high-resolution magnetometry Tschirhart et al. (2021); Grover et al. (2022).
@@
Transport and scanning tunneling microscopy experiments further showed that weak to moderate out-of-plane magnetic fields can drive the states at $\nu=\pm 1,\pm 2,\pm 3$ into correlation-induced Chern insulators in magic-angle TBG Das et al. (2021); Nuckolls et al. (2020). These observations highlight orbital magnetization as a key quantity governing the coupling of correlated moiré states to external magnetic fields. A quantitative understanding of orbital magnetization is therefore essential for elucidating the microscopic origin of Chern-insulating phases and field-induced transitions between competing states in TBG Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022).
@@
On the theoretical side, the Berry curvature and orbital magnetic moments of moiré bands have been extensively analyzed within the non-interacting Bistritzer–MacDonald (BM) continuum model Liu et al. (2019b); He et al. (2020); Liu and Dai (2021b); Zhu et al. (2020). These studies established that the flat bands of TBG can host nontrivial topology and substantial orbital magnetization per unit cell largely arising from sizable Berry curvature. However, the flat-band nature of magic-angle TBG implies that electron–electron interactions are comparable to or even larger than the bandwidth, and self-consistent Hartree–Fock (HF) calculations have demonstrated that interactions tremendously reconstruct the band structure and stabilize correlated Chern insulating states at integer fillings Kang and Vafek (2019, 2020); Zhang et al. (2019, 2020); Xu et al. (2018); Xie and MacDonald (2020); Liu et al. (2018); Huang et al. (2019); Liu and Dai (2021a); Liu et al. (2021); Bultinck et al. (2020a, b); Wu and Das Sarma (2020); Chen et al. (2021); Lu et al. (2022); Da Liao et al. (2021); Bernevig et al. (2021); Lian et al. (2021); Xie et al. (2021b); Hejazi et al. (2021); Sboychakov et al. (2019); Rademaker et al. (2019); Guinea and Walet (2018); Seo et al. (2019); Kang et al. (2021); Wang and Vafek (2024, 2022). Despite this progress, a systematic calculation of the bulk orbital magnetization in such interaction-driven states remains lacking. In particular, it is not clear how interaction-induced band reconstruction, valley polarization, and remote-band contributions quantitatively modify the orbital magnetization beyond the non-interacting predictions, nor how these effects relate to experimentally observed field-induced switching between competing correlated phases.
In this work, we combine the recently developed projector formulation of orbital magnetization Kang et al. (2025); Zhu and Huang (2025); Liu et al. (2025); Ye (2026); Chen and Song (2026) with self-consistent Hartree-Fock theory to construct a gauge-invariant framework for correlated states of magic-angle TBG Ceresoli et al. (2006). Starting from the BM continuum model with Coulomb interactions projected onto the relevant bands, we determine symmetry-broken ground states at integer fillings and evaluate the orbital magnetization using the modern theory expressed in terms of projection operators. This formulation avoids gauge ambiguities and enables a controlled treatment of the large number of remote bands inherent to the continuum description.
@@
Using this framework, we compute both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ relevant for magneto-optical probes such as magnetic circular dichroism (MCD) Xia et al. (2026). We show that, unlike topological invariants such as the Chern number, the orbital magnetization is strongly sensitive to remote-band contributions and therefore requires careful convergence. The calculated orbital magnetization respects the particle-hole symmetry that relates the fillings of $\nu$ and $\nu$ Bernevig et al. (2021) and reproduces the characteristic linear dependence of magnetization on chemical potential within the insulating gap, with slope proportional to the Chern number. We further analyze the sign structure of the magnetization and its relation to changes in Hall conductivity. At $\nu=\pm 1$, we compare competing Chern and intervalley-coherent states and show that Chern insulating phases possess substantially larger orbital magnetization, providing a microscopic explanation for their stabilization under external magnetic fields. Our results establish a systematic approach for evaluating orbital magnetization in correlated moiré systems and elucidate the important role of interactions as well as the critical contribution of remote bands in shaping orbital ferromagnetism in twisted bilayer graphene.
@@ II Non-interacting model
where $\hat{{\mathbf{k}}}=-i\partial_{{\mathbf{r}}}$, $\tau=\pm$ is the valley index, and $\boldsymbol{\sigma}^{\tau}=(\tau\sigma_{x},\sigma_{y})$ the Pauli matrices acting in the A,B sublattice space. In Eq. (1), ${\mathbf{K}}^{\tau}_{1}$ and ${\mathbf{K}}^{\tau}_{2}$ are the Dirac points of valley $\tau$ in the bottom and top layers twisted by angles $\mp\frac{\theta}{2}$. They are equivalent to $\tau{\mathbf{K}}_{1}$ and $\tau{\mathbf{K}}_{2}$ in the moiré Brillouin zone (mBZ), as shown in Fig. 1(c). The interlayer tunneling between the Dirac states in the two layers is described by the matrix
@@
where $u_{0}$ and $u_{1}$ are the intra- and inter-sublattice interlayer tunneling amplitudes. We adopt realistic parameters for the continuum model: $\hbar v_{F}/a_{0}=2.365$ eV, inter-sublattice interlayer coupling $u_{1}=0.11$ eV, and intra-sublattice interlayer coupling $u_{0}=0.06$ eV. The twist angle is fixed at the magic value $\theta$=1.086, at which the low-energy moiré bands become perfectly flat in the limit $u_{0}\to 0$ Tarnopolsky et al. (2019). The low-energy flat bands exhibit a bandwidth of approximately 2 meV and are well separated from the higher-energy remote bands by an energy gap of about 50 meV as shown in Fig. 1. The eigenstates of the continuum Hamiltonian $H_{BM}^{\tau}$ have the Bloch wavefunction form
@@
where $X=\{A_{1},B_{1},A_{2},B_{2}\}$ is the layer and sublattice index with the eigen-band energy $\epsilon_{m{\mathbf{k}}\tau}$ and $\mathbf{G}$} is the general reciprocal lattice vector in the mBZ. Here, $m$ and $\tau$ are the band and valley indices, and the spin index $s$ is omitted since the Hamiltonian is spin independent. In the calculation, we take 121 $\mathbf{G}$} vectors as expressed as ${\mathbf{G}}=n_{1}{\mathbf{G}}_{1}+n_{2}{\mathbf{G}}_{2}$ with $n_{1},n_{2}\in[-5,5]$, which is enough to produce the flat bands.
Figure 1: (a) Low-energy band structure of non-interacting TBG calculated using the BM model. The black solid and red dashed curves correspond to the $K^{\prime}$ and $K$ valleys, respectively. (b) Zoomed-in view of the isolated flat bands near charge neutrality, highlighting their separation from the higher-energy remote bands. (c) Schematic illustration of Brillouin zone folding in TBG. The small hexagon denotes the moiré Brillouin zone (mBZ), reciprocal to the moiré superlattice. The cyan arrows indicate the momentum path in the enlarged moiré Brillouin zone along which the band dispersions in panels (a) and (b) are plotted.
@@ III Interacting case
The full Hamiltonian is written as $H=H_{0}+H_{int}$, with the non-interacting part given by
$H_{0}=\sum_{\tau}H_{BM}^{\tau}=\sum_{m{\mathbf{k}}s\tau}\epsilon_{m{\mathbf{k}}\tau}d_{{\mathbf{k}},m,s,\tau}^{\dagger}d_{{\mathbf{k}},m,s,\tau},$ (4)
Here, $d^{\dagger}_{{\mathbf{k}},m,s,\tau}$ and $d_{{\mathbf{k}},m,s,\tau}$ are the creation and annihilation operators in the eigen-band basis, respectively. To incorporate interactions, we project the single-gate-screened Coulomb interaction onto the two lowest-energy flat bands for each spin and valley flavor, and solve the resulting problem self-consistently within a standard Hartree-Fock (HF) approximation in momentum space. Within this treatment, interaction effects are retained only in the active flat-band subspace, while the remote bands are kept fixed. The detailed construction of the projected interaction, the form factors, and the resulting HF Hamiltonian is presented in the Supplemental Material sm , following the standard formulations of Refs. Bernevig et al. (2021); Kang et al. (2021); Zhang et al. (2020). In the numerical calculations, we adopt a dielectric constant $\varepsilon=7$ and a gate distance $d_{s}=40$ nm for the single gate screened Coulomb interaction. In the HF calculation, we take $N_{k}=30\times 30=900$ kpoints in the mBZ.
@@
It is well established that electron–electron Coulomb interactions in twisted bilayer graphene can drive spontaneous symmetry breaking, giving rise to a series of correlated insulating states at commensurate integer fillings. Among these, states that break the combined twofold-rotation and time-reversal symmetry $C_{2}\mathcal{T}$, where $\mathcal{T}$ denotes time reversal, are energetically competitive. In particular, at fillings $\nu=\pm 3$, the TRSB Chern insulating states can emerge as ground states even in the absence of an explicit staggered sublattice potential. Within this projected-basis HF framework, we systematically identify the correlated insulating states at all integer fillings and compute the orbital magnetization of these interaction-driven phases.